# Learning Prompt-Enhanced Context Features for Weakly-Supervised Video Anomaly Detection

### v1

@@ IV-E Analysis of Time–Space Complexity
(d) Normal_Videos_783_x264
(e) Normal_Videos_887_x264
(f) Normal_Videos_725_x264
Figure 10: Qualitative results of our method on the UCF-Crime dataset. The green and red boxes represent normal and abnormal keyframes, respectively. The second and third rows indicate the results of ‘w/ TCA’ and ‘w/ PEL & SS’, respectively. (a)-(e) are correct videos and (f) is a false alarm sample.
(a) Ip.Man.2008
(b) The.Fast.and.the.Furious.2001
(c) v=0yHBkMBE8r4
(d) Bullet.in.the.Head.1990
(e) City.Of.Men.2007
(f) v=lVJVRywgmYM
Figure 11: Qualitative results of our method on the XD-Violence dataset. The green and red boxes represent normal and abnormal keyframes, respectively. The second and third rows indicate the results of ‘w/ TCA’ and ‘w/ PEL & SS’, respectively. (a)-(e) are correct videos and (f) is a false alarm sample.
IV-F Qualitative Results
We first visualize the correlation of context features to demonstrate the encoding performance of the TCA module. As shown in Figure 8, the heatmap demonstrates the cosine similarity between the snippets, where the first two rows are test videos of UCF-Crime and the last two rows are from XD-Violence. We observe that 1) the variability of UCF-Crime videos is less pronounced than that of XD-Violence since the former is captured from fixed surveillance while the latter contains complex scene and camera changes; 2) local features are more discriminable than global features because the latter introduces more long-range noise; and 3) local-global aggregated features can effectively suppress redundancies and further enhance the discriminability among snippets.
To further illustrate the efficacy of the PEL module, we utilize t-SNE to visualize the features of the middle layer of the MLP. As depicted in Figure 9, the distribution of test samples in the UCF-Crime dataset appears disorderly prior to alignment, whereas the samples in the XD-Violence dataset exhibit a discernible clustering pattern. We contend that the former surveillance videos possess temporal coherence, resulting in blurred fore-background boundaries. Conversely, the latter videos have natural shot boundaries created by artificial means, leading to a more distinguishable boundary. Through the incorporation of cross-modal alignment loss, abnormal foregrounds converge towards the corresponding prompt features to form compact clusters. In contrast, hard-negative snippets and backgrounds align with a uniform normal center, consequently increasing the relative distance between abnormal and normal snippets in the embedding space. Noteworthy is that distinct fine-grained anomalous samples are effectively differentiated, surpassing current binary classification models in terms of semantic representation.
Finally, we report the anomaly scores of several test videos in Figures 10 and 11. The second row displays the detection results of the TCA module, while the third row shows the outcomes of PEL and SS. Compared to TCA, the introduction of PEL and SS can achieve more accurate anomaly localization results, which can be attributed to the effective suppression of non-anomaly noise through F-B separation in PEL. Meanwhile, false alarms triggered by in/out-of-frames and luminance changes are corrected, thereby further reducing the false alarm rate in normal videos. Our model can be adapted to multiple scenarios and types of anomalous event detection, owing to the rich semantic information provided by the prompt for context features.
@@ V Conclusion
In this paper, we focus on efficient temporal context modeling and semantic enhancement of visual features for weakly supervised video anomaly detection. Our model achieves local-global context aggregation through the reuse of the similarity matrix and introduces knowledge-based prompts to enhance the weak semantic discriminability of context features. The incorporation of context information and external knowledge significantly improves the accuracy of anomaly localization. In addition, we implemented score smoothing to reduce individual bias in continuous sequences and reduce false alarms. The results of extensive experiments demonstrate that our model performs competitively on three benchmarks with fewer parameters and computational cost. Additionally, the detection accuracy of several abnormal sub- classes is significantly improved. In the future, multimodal information such as motion and audio remains to be explored, and open-set anomaly detection also deserves attention.
Acknowledgments
This work was supported by the state key development program in the 14th Five-Year under Grant No.2021YFF0900701, 2021YFF0602103, 2021YFF0602102, 2021QY1702, and in part by National Natural Science Foundation of China (No.61801441). We also thank the research funds under Grant No.2019GQG0001 from the Institute for Guo Qiang, Tsinghua University, and the High-quality and Cutting-edge Disciplines Construction Project for Universities in Beijing (Internet Information, Communication University of China).

### v2

@@ IV-D6 Impact of Score Smoothing
(d) Ip.Man.2008
(e) The.Fast.and.the.Furious.2001
(f) v=0yHBkMBE8r4
Figure 8: Anomaly score visualization of the proposed method. Orange regions indicate ground-truths and blue curves indicate anomaly scores. (a)-(c) are videos from UCF-Crime and (d)-(f) are from XD-Violence, where the second and third rows indicate the results of ‘w/ TCA’ and ‘w/ PEL & SS’, respectively.
(a) Assault010_x264
(b) Explosion027_x264
(c) Fighting003_x264
Figure 9: Anomaly score visualization of different methods on the UCF-Crime dataset. The first row is from our proposed framework, and the second and third rows show the detection results of DDL[55] and CUPL[15], respectively.
IV-E Parameter Evaluation
In this subsection, we explore the effect of different hyperparameter settings on model performance.
IV-E1 Local Window Size in TCA module
Figure 6(a) demonstrates the effect of different local window sizes in the TCA module. The performance of the model rises and then falls on both datasets as the window continues to increase, with the best results achieved on both datasets when the window $w$ is set to 9. In general, smaller windows lack sufficient contextual information, while larger windows introduce long-range noise, leading to oscillations in false alarm rates.
IV-E2 Kernel Size $\Delta t$ in Classifier
Figure 6(b) illustrates the impact of varying causal convolution kernel sizes on the UCF-Crime dataset. It is observed that the best results are achieved with a kernel size of 9. However, at this kernel size, the false alarm rate is not optimal, indicating a higher number of normal samples being misclassified. This suggests a need for further refinement in the model to balance detection accuracy and false alarm rates effectively.
IV-E3 Pooling Size $\kappa$ for Score Smoothing
We also investigate the effect of different pooling window sizes in the SS strategy, as presented in Figure 6 (c) and (d). The results show that our model is more sensitive to the size of the sliding window compared to the moving window. The optimal performance on the UCF-Crime dataset is achieved when the sliding window size is set to 7, resulting in a false alarm rate of 0.43%. For the XD-Violence dataset, a moving window size of 9 works best, with a further reduction in the FAR to 0.57%.
TABLE XIII: Performance comparison of different temperature coefficients $\tau$
$\tau$ 0.01 0.03 0.05 0.07 0.09 0.1
UCF@AUC (%) 85.49 85.94 85.84 85.61 86.36 85.98
XD@AP (%) 84.11 83.56 85.26 84.16 84 84.63
TABLE XIV: Performance comparison of different loss weight $\lambda$
$\lambda$ 0.01 0.1 0.5 1 5 10
UCF@AUC (%) 85.11 85.19 85.52 86.36 84.51 84.42
XD@AP (%) 82.77 83.3 84.48 85.26 82.65 83.43
IV-E4 Temperature Coefficient $\tau$ and Loss Weight $\lambda$
Tables XIII and XIV report the model’s performance for different loss weights and temperature coefficients. Our model achieves optimal performance on both datasets when the weight is set to 1, which ensures a good balance of classification and alignment terms. However, the temperature coefficient varies for each dataset, with the best performance on UCF-Crime achieved at 0.09 and on XD-Violence at 0.05.
IV-F Qualitative Analysis
We first use t-SNE visualization for the MLP’s middle layer features, as shown in Figure 7. Before introducing PEL, the UCF-Crime dataset samples appear disordered, while the XD-Violence dataset shows clearer clustering. This difference is attributed to the temporal coherence in surveillance videos and distinct shot boundaries in XD-Violence content. Cross-modal alignment loss makes abnormal foregrounds converge towards prompt features, forming tighter clusters. It also aligns hard-negative snippets and backgrounds with a normal center, increasing the distance between abnormal and normal snippets in the embedding space. This results in a more effective differentiation of fine-grained anomalous samples compared to traditional binary classification models.
Figure 8 illustrates the anomaly scores obtained using our methods. The TCA module’s detection results are shown in the second row, while the combined effect of PEL and SS is presented in the third row. PEL and SS, when used together, achieve more precise anomaly localization than TCA alone. This increased accuracy is attributed to the effective suppression of non-anomaly noise by PEL’s context separation and the mitigation of false alarms caused by frame changes and luminance variations. Moreover, the integration of rich semantic information from prompts significantly enhances the model’s versatility in detecting various anomalous events across different scenarios.
Finally, we compare the detection effectiveness of various methods in Figure 9. The PEL method exhibits superior sensitivity and specificity in identifying anomalies within the UCF-Crime dataset, particularly evident in scenarios like Assault, Explosion, and Fighting. PEL’s anomaly scoring is characterized by sharp peaks and clear distinctions of anomalous events, in stark contrast to the broader, less defined detection of DDL and the delayed or dispersed responses of CUPL. This demonstrates PEL’s exceptional temporal accuracy and event localization capabilities, establishing it as a highly effective tool for real-world surveillance applications.
@@ V Conclusion
In this paper, we focus on efficient temporal context modeling and semantic enhancement of visual features for weakly supervised video anomaly detection. We introduce a temporal context aggregation module that reuses the similarity matrix to capture local-global dependencies simultaneously. This method not only cuts down on parameters and computational load but also boosts detection capabilities, marking a significant shift from conventional paralleled systems and enhancing real-world applicability and efficiency. Recognizing the complex nature of anomalies, our model doesn’t just detect but understand them by integrating external knowledge for semantic interpretability. This aligns the model with human cognition, offering nuanced understanding be yond typic al binary constraints. Our proposed prompt-enhanced learning module further refines this by using class-specific prompts derived from external knowledge, improving the distinction between abnormal sub-classes and maintaining clear inter-class separation. In the future, multimodal information such as motion and audio remains to be explored, and open-set anomaly detection also deserves attention.